### v1

@@ 6 Conclusion
In this paper, we introduced Pangea, a novel multilingual multimodal large language model designed to bridge linguistic and cultural gaps in visual understanding tasks. By leveraging PangeaIns, our newly curated 6M multilingual multimodal instruction data samples, we demonstrated significant improvements in cross-lingual and cross-cultural understanding across 39 typologically diverse languages. Our comprehensive evaluation using PangeaBench revealed Pangea’s superior performance compared to existing open-source models, particularly in tasks requiring nuanced cultural interpretation. We also highlight ongoing challenges in areas such as low-resource language support and multilingual OCR. We fully open-source Pangea, PangeaIns, and PangeaBench to facilitate future research to build open and inclusive MLLMs.
Acknowledgments
This work was supported in part by the Carnegie Bosch Institute Fellowship and a grant from DSTA Singapore. The training is supported by the CMU FLAME Center. The authors would like to thank CMU NeuLab colleagues for their constructive comments. The authors would like to thank Google Gemini credits for data construction and evaluation.
@@ Appendix A Related Work
Visual instruction tuning is a key technique for enhancing multimodal large language models by aligning visual inputs with textual instructions to improve understanding and generation tasks Liu et al. (2023b). Traditionally, these instructions are built using English-language data from visual question answering and other datasets Liu et al. (2023b; a; 2024); Tong et al. (2024); Beyer et al. (2024). Researchers often supplement this with synthetic instruction tuning data, generating large volumes of instructional pairs to possibly cover multiple languages too Geigle et al. (2024). However, these instruction-tuning datasets have mostly been task-focused and lack conversational capabilities. Further, while translation gives lends to multilingual capabilities, the data remains to be culturally homogeneous. By curating multilingual and multicultural instruction tuning data across various task types, our model is designed to intuitively understand and engage with users from diverse demographics.
@@
Multilingual MLLMs have evolved from dual-encoder-based models, only capable of understanding and reasoning Ni et al. (2021); Zeng et al. (2022); Jain et al. (2021), to encoder-decoder models capable of multilingual text generation as well Shan et al. (2022); Chen et al. (2023); Geigle et al. (2024). Despite their advancements, these models have remained focused on conventional tasks such as VQA and image captioning. Moreover, most efforts have centered around training with multilingual text, while little attention has been given to curating culturally diverse image datasets. Even for text, despite the focus on multilinguality, few attempts have been made to reflect cultural diversity in instructions and captions. As a result, these models tend to reflect a Western-centric bias. By selecting culturally diverse images from LAION and intentionally integrating this diversity into our instructions and captions, our model aims to serve a wide range of users in an inclusive and equitable manner.
@@ Appendix E Explanation of xChatBench
Similar to prior works employing LLM-as-a-Judge, we use GPT-4o-2024-08-06 as the judge model and prompt it in a direct assessment manner. As input, the judge model is given the instruction, the model’s response, the reference answer, the evaluation criteria, and the descriptions for each score. As output, the judge generates verbal feedback and an integer score between 1 and 5. For this procedure, we use the prometheus-eval library (Kim et al., 2024b) and employ their default hyperparameter setting for evaluation. Lastly, the final score is acquired by averaging the results across the 50 instances for each language. Note that in the main result and breakdown result tables, we normalize the score from 1-5 to 0-100 by $(\text{score}-1)\times 25$. Also, for postprocessing, we use langdetect  [^3] https://pypi.org/project/langdetect/ to identify whether the response is written in the given language and change the score to 1 when it is written in a different language, a phenomenon called language hallucination (Xue et al., 2021; Pfeiffer et al., 2023).
@@ Appendix F Qualitative Examples from xChatBench
Appendix G Training Examples
G.1 Machine Translated Instructions
@@
G.2 Multicultural Understanding Instructions
@@
Appendix H Breakdown Results of Different Languages on PangeaBench
H.1 xChat
We show the performance of different models’ performance on the xChat benchmark in Table 3.
@@ Appendix H Breakdown Results of Different Languages on PangeaBench
H.2 Multilingual LLaVABench
We show the performance of different models’ performance on the Multilingual LLaVABench benchmark in Table 4.
@@ H.2 Multilingual LLaVABench
H.3 CVQA
We show the performance of different models’ performance on the CVQA benchmark in Table 5 and Table 6.
@@ H.3 CVQA
H.4 MaRVL
We show the performance of different models’ performance on the MaRVL benchmark in Table 7.
@@ H.4 MaRVL
H.5 XM100
We show the performance of different models’ performance on the XM100 benchmark in Table 8.
@@ H.5 XM100
H.6 xGQA
We show the performance of different models’ performance on the xGQA benchmark in Table 9.
@@
H.7 MAXM
We show the performance of different models’ performance on the MAXM benchmark in Table 10.
@@ H.7 MAXM
H.8 xMMMU
We show the performance of different models’ performance on the xMMMU benchmark in Table 11.
@@
H.9 M3Exam
We show the performance of different models’ performance on the M3Exam benchmark in Table 12.
@@ H.9 M3Exam
H.10 TyDiQA
We show the performance of different models’ performance on the TyDiQA benchmark in Table 13.
@@ H.10 TyDiQA
H.11 XStoryCloze
We show the performance of different models’ performance on the XStoryCloze benchmark in Table 14.
@@ H.11 XStoryCloze
H.12 MGSM
We show the performance of different models’ performance on the MGSM benchmark in Table 15.
@@ H.12 MGSM
H.13 MMMLU
We show the performance of different models’ performance on the MMMLU benchmark in Table 16.
@@ H.13 MMMLU
Appendix I A Preliminary Exploration of Constructing Multilingual OCR Instructions
Optical Character Recognition (OCR) is a critical capability for multimodal LLMs, enabling them to interpret and process textual information embedded within images. However, most existing OCR training datasets are predominantly English-centric, which limits the models’ performance in non-English contexts. To address this gap, we have curated a comprehensive set of 500K multilingual OCR training samples from web user interfaces.

### v2

@@ 6 Conclusion
In this paper, we introduced Pangea, a multilingual MLLM designed to bridge linguistic and cultural gaps in visual understanding tasks. By leveraging PangeaIns, our newly curated 6M multilingual multimodal instruction data samples, we demonstrated significant improvements in cross-lingual and cross-cultural understanding across 39 typologically diverse languages. Our comprehensive evaluation using PangeaBench revealed Pangea’s superior performance compared to existing open-source models. We also highlight ongoing challenges in areas such as low-resource language support and multilingual OCR. We fully open-source Pangea, PangeaIns, and PangeaBench to facilitate future research to build open and inclusive MLLMs.
@@ Appendix A Related Work
Visual instruction tuning is a key technique for enhancing multimodal large language models by aligning visual inputs with textual instructions to improve understanding and generation tasks Liu et al. (2023b). Traditionally, these instructions are built using English-language data from visual question answering and other datasets Liu et al. (2023b; a); Xu et al. (2023); Liu et al. (2024); Tong et al. (2024); Beyer et al. (2024); Zhu et al. (2024); Dai et al. (2023); Tong et al. (2024); Li et al. (2024d). Researchers often supplement this with synthetic instruction tuning data, generating large volumes of instructional pairs to possibly cover multiple languages too Geigle et al. (2024b); Li et al. (2023; 2024d). However, these instruction-tuning datasets have mostly been task-focused and lack conversational capabilities. Further, while translation gives lends to multilingual capabilities, the data remains to be culturally homogeneous. By curating multilingual and multicultural instruction tuning data across various task types, our model is designed to intuitively understand and engage with users from diverse demographics.
Dataset # Languages # of Instances Multicultural # of Task Types Open-Sourced
MultiInstruct (Xu et al., 2023) 1 $\sim$235.0K 310
MiniGPT4 (Zhu et al., 2024) 1 5.0K 149 $\checkmark$
LLAVA (Liu et al., 2023b) 1 1.2M $>$100K $\checkmark$
InstructBLIP (Dai et al., 2023) 1 $\sim$1.6M $>$100K
M3IT (Li et al., 2023) 80 2.4M 400 $\checkmark$
mBLIP (Geigle et al., 2024a) 95 5.1M 68 $\checkmark$
PALO (Rasheed et al., 2025) 10 2.1M 22 $\checkmark$
Cambrian (Tong et al., 2024) 1 7.1M $>$1M $\checkmark$
PangeaIns (Ours) 39 6.2M $\checkmark$ $>$1M $\checkmark$
Table 4: Comparison of datasets in terms of number of languages, number of instances, whether the dataset is multicultural, number of task types, and open-sourced.
@@
Multilingual MLLMs have evolved from dual-encoder-based models, only capable of understanding and reasoning Ni et al. (2021); Zeng et al. (2022); Jain et al. (2021), to encoder-decoder models capable of multilingual text generation as well Shan et al. (2022); Chen et al. (2023); Geigle et al. (2024b). Despite their advancements, these models have remained focused on conventional tasks such as VQA and image captioning. Moreover, most efforts have centered around training with multilingual text, while little attention has been given to curating culturally diverse image datasets. Even for text, despite the focus on multilinguality, few attempts have been made to reflect cultural diversity in instructions and captions. As a result, these models tend to reflect a Western-centric bias. By selecting culturally diverse images from LAION and intentionally integrating this diversity into our instructions and captions, our model aims to serve a wide range of users in an inclusive and equitable manner.
@@ Appendix E Explanation of xChatBench
Similar to prior works employing LLM-as-a-Judge, we use GPT-4o-2024-08-06 as the judge model and prompt it in a direct assessment manner. As input, the judge model is given the instruction, the model’s response, the reference answer, the evaluation criteria, and the descriptions for each score. As output, the judge generates verbal feedback and an integer score between 1 and 5. For this procedure, we use the prometheus-eval library (Kim et al., 2024b) and employ their default hyperparameter setting for evaluation. Lastly, the final score is acquired by averaging the results across the 50 instances for each language. Note that in the main result and breakdown result tables, we normalize the score from 1-5 to 0-100 by $(\text{score}-1)\times 25$. For the multimodal chat scenarios, we found that many English-centric models tend to respond in English regardless of the query language. This behavior is problematic, as it undermines the fundamental capability of a multilingual model, which should ideally respond in the language of the query. To address this, we implemented a strict evaluation criterion where such responses were penalized and assigned a score of 0. We believe this is crucial, as users may not understand English, and failing to respond in the appropriate language can hinder effective communication and user experience. Thus, for postprocessing, we use langdetect  [^2] https://pypi.org/project/langdetect/ to identify whether the response is written in the given language and change the score to 1 when it is written in a different language, a phenomenon called language hallucination (Xue et al., 2021; Pfeiffer et al., 2023).
@@ Appendix F Qualitative Examples from xChatBench
Appendix G Languages in PangeaIns
Table 5 demonstrates the number of instances for each language that we include in PangeaIns.
Languages en multi am ar bg bn cs de el es
Count 2554.6 4389.5 31.7 162.8 52.7 118.4 4.6 132.2 7.3 126.6
Percentage (%) 36.8 63.2 0.5 2.3 0.8 1.7 0.1 1.9 0.1 1.8
Languages fa fr ga hi id ig it iw ja jv
Count 8.4 193.8 34.7 163.5 178.4 26.1 68.8 168.3 403.3 35.2
Percentage (%) 0.1 2.8 0.5 2.4 2.6 0.4 1.0 2.4 5.8 0.5
Languages ko nl mn ms no pl pt ro ru si
Count 182.5 4.5 37.6 39.2 60.7 8.2 176.5 147.2 211.0 0.6
Percentage (%) 2.6 0.1 0.5 0.6 0.9 0.1 2.5 2.1 3.0 0.1
Languages su sw ta te th tr uk ur vi zh
Count 33.0 132.7 37.1 64.6 160.4 140.7 7.4 156.4 315.5 557.4
Percentage (%) 0.5 1.9 0.5 0.9 2.3 2.0 0.1 2.3 4.5 8.0
Table 5: Language distribution of PangeaIns. We demonstrate the number of instances (in thousands) for each language, and the percentage of instances for each language in PangeaIns
Appendix H Training Examples
H.1 Machine Translated Instructions
@@
H.2 Multicultural Understanding Instructions
@@
Appendix I Breakdown Results of Different Languages on PangeaBench
I.1 xChat
We show the performance of different models on the xChat benchmark in Table 6.
@@ Appendix I Breakdown Results of Different Languages on PangeaBench
I.2 Multilingual LLaVABench
We show the performance of different models on the Multilingual LLaVABench benchmark in Table 7.
@@ I.2 Multilingual LLaVABench
I.3 CVQA
We show the performance of different models on the CVQA benchmark in Table 8 and Table 9.
@@ I.3 CVQA
I.4 MaRVL
We show the performance of different models on the MaRVL benchmark in Table 10.
@@ I.4 MaRVL
I.5 XM100
We show the performance of different models on the XM100 benchmark in Table 11.
@@ I.5 XM100
I.6 xGQA
We show the performance of different models on the xGQA benchmark in Table 12.
@@
I.7 MAXM
We show the performance of different models on the MAXM benchmark in Table 13.
@@ I.7 MAXM
I.8 xMMMU
We show the performance of different models on the xMMMU benchmark in Table 14.
@@
I.9 M3Exam
We show the performance of different models on the M3Exam benchmark in Table 15.
@@ I.9 M3Exam
I.10 TyDiQA
We show the performance of different models on the TyDiQA benchmark in Table 16.
@@ I.10 TyDiQA
I.11 XStoryCloze
We show the performance of different models on the XStoryCloze benchmark in Table 17.
@@ I.11 XStoryCloze
I.12 MGSM
We show the performance of different models on the MGSM benchmark in Table 18.
@@ I.12 MGSM
I.13 MMMLU
We show the performance of different models on the MMMLU benchmark in Table 19.
@@ I.13 MMMLU
Appendix J A Preliminary Exploration of Constructing Multilingual OCR Instructions
Optical Character Recognition (OCR) is a critical capability for multimodal LLMs, enabling them to interpret and process textual information embedded within images. However, most existing OCR training datasets are predominantly English-centric, which limits the models’ performance in non-English contexts. To address this gap, we have curated a comprehensive set of 500K multilingual OCR training samples from web user interfaces, spanning 10 languages, with 50K examples per language, sourced from web user interfaces. Webpages naturally serve as image-rich environments containing abundant text, and by capturing screenshots of websites from various countries in different languages, we were able to gather a substantial number of OCR images.